OpenStudy (anonymous):

Find the exact volume of the solid between the paraboloids z=2x^2+y^2 and z=8−x^2−2y^2 and inside the cylinder x^2+y^2=1

OpenStudy (anonymous):

I really don't know how to set this up. Would it be something like ∫∫(2x^2+y^2)-(8-x^2-2y^2)dA + ∫∫(x^2+y^2-1)dA ? If so, how would I find the bounds of integration?

OpenStudy (anonymous):

int tita=0-2pi r=0-1 z=r2-8-r2

OpenStudy (anonymous):

What are you doing? Not sure I understand.

OpenStudy (anonymous):

rdzdrdtita =8r-2r3 dr=4r2-1/2r4]0-1=3.5 3.5(2pi)=7pi

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

unfortunetly my bamboo not work properly to show you. but I calculated triple integral.

OpenStudy (anonymous):

How? How would you do this with a double integral?

OpenStudy (anonymous):